sunshine clinic purchases a piece of equipment for $250 with a 2-15 net provision. what is the effective interest rate if the clinic pays on day 16? on day 30? a. 292 percent on day 16; 19.5 percent on day 30 b. 146 percent on day 16; 9.7 percent on day 30 c. 146 percent on day 16; 19.5 in day 30 d. 292 percent on day 16; 9.7 percent on day 30.
Answer #1

The options are not right- please recheck your options.. the correct answer is 745 percent on day 16; 49.7 percent on day 30

Solution-

2% off in 15 day
16-15 = 1
365/1 = 365

250*0.02 = 5

365*5 = 1825

250 * .98 = 245

1825/245 = 7.44898 = 7..448*100=744.8=745%

for day 30

30-15 = 15

(365/15) * 5 = 121.66667

250 * .98 = 245

121.66667/245 =0.4965 =0.4965*100= 49.6 %